# TEMPO: Temporal Preference Optimization of Video LLMs via Difficulty Scheduling and Pre-SFT Alignment

### v1

@@ Abstract
Video Large Language Models (Video LLMs) have achieved significant success by leveraging a two-stage paradigm: pretraining on large-scale video-text data for vision-language alignment, followed by supervised fine-tuning (SFT) for task-specific capabilities. However, existing approaches struggle with temporal reasoning due to weak temporal correspondence in the data and reliance on the next-token prediction paradigm during training. To address these limitations, we propose TEMPO (TEMporal Preference Optimization), a systematic framework that enhances Video LLMs’ temporal reasoning capabilities through Direct Preference Optimization (DPO). To facilitate this, we introduce an automated preference data generation pipeline that systematically constructs preference pairs by selecting videos that are rich in temporal information, designing video-specific perturbation strategies, and finally evaluating model responses on clean and perturbed video inputs. Our temporal alignment features two key innovations: curriculum learning which that progressively increases perturbation difficulty to improve model robustness and adaptability; and “Pre-SFT Alignment”, applying preference optimization before instruction tuning to prioritize fine-grained temporal comprehension. Extensive experiments demonstrate that our approach consistently improves Video LLM performance across multiple benchmarks with a relatively small set of self-generated DPO data. We further analyze the transferability of DPO data across architectures and the role of difficulty scheduling in optimization. Our findings highlight our TEMPO as a scalable and efficient complement to SFT-based methods, paving the way for developing reliable Video LLMs.
@@ 1 Introduction
To address these limitations, in this paper, we propose TEMPO (TEMporal Preference Optimization), a systematic framework that tackles both challenges through: (1) a scalable, automatic dataset curation pipeline that generates high-quality temporal preference data, and (2) a tailored temporal alignment methodology that goes beyond next-token prediction. TEMPO is built on Direct Preference Optimization (DPO) [24], which optimizes the relative quality of model responses by maximizing the reward margin between preferred and non-preferred outputs. Unlike SFT, which requires high-quality, human-verified responses, DPO focuses on relative comparisons rather than absolute correctness—making it particularly effective for video tasks where perfect annotations are difficult to obtain. Subtle temporal aspects like sequence order and motion speed are critical for comprehensive video understanding yet challenging to annotate manually at scale. Our preference-based framework naturally discourages shortcut learning while enabling efficient scaling through automated annotation, making TEMPO an ideal methodology for enhancing temporal reasoning in Video LLMs.
@@
We implement TEMPO through an automatic preference data generation pipeline that establishes the foundation for effective temporal alignment in Video LLMs. Inspired by advances in image-based LLMs [33, 41], we generate preference pairs by evaluating model responses under both clean and corrupted visual inputs. Our pipeline consists of three key stages: (1) curating a diverse set of videos that emphasize complex temporal relationships; (2) introducing targeted temporal perturbations—including frame dropping, sequence reversal, and frame shuffling—that precisely target known temporal reasoning weaknesses of current Video LLMs; and (3) constructing preference pairs by contrasting the model’s outputs for clean versus temporally corrupted videos. We also introduce a difficulty factor that quantifies perturbation magnitude for controlling the sample difficulties. This systematic approach enables us to efficiently generate high-quality contrastive examples that directly highlight temporal understanding capabilities without requiring expensive human annotation.
@@
Our tailored temporal alignment framework consists of two key technical innovations. First, we implement curriculum learning with progressive difficulty lifting, which gradually increases sample complexity during optimization. This controlled exposure to increasingly challenging temporal constrastive pairs enables more effective learning. Second, we introduce a novel “Pre-SFT Alignment” strategy, which reverses the conventional order of model alignment. Traditional approaches apply DPO after instruction-tuning (SFT) [30], assuming the model has already learned basic capabilities. In contrast, we argue that robust temporal understanding should be established first, before learning to follow diverse instructions. Our Pre-SFT Alignment applies DPO first on our curated dataset, forcing the model to develop strong temporal reasoning and video-text correspondence prior to any instruction tuning.
@@
Through extensive evaluation, we demonstrate that TEMPO consistently improves model performance across multiple benchmarks including Video-MME [8], MLVU [40] and Vinoground [36] with only a small number of self-generated DPO pairs—eliminating dependence on external models or APIs. Our ablation studies validate the effectiveness of both technical contributions: the difficulty curriculum significantly stabilizes training and improves final performance, while the Pre-SFT Alignment strategy significantly outperforms the conventional SFT-then-DPO approach, particularly on tasks requiring fine-grained temporal reasoning. Further analysis reveals valuable insights about DPO data transferability between different model architectures and scales, demonstrating the framework’s flexibility. By introducing this automated, structured preference learning pipeline, we advance the state of video LLMs and pave the way for more robust and temporally-aware multi-modal models.
@@ 3 Experiment
   + TEMPO (Ours) 7B 74.5  (+3.6) 45.2  (+2.8) 61.0  (+0.6) 65.8  (+1.3) 37.2  (+3.2) 29.6  (+3.0) 12.2  (+1.6) 26.3  (+2.6)
Table 1: Benchmark results of our approach on VideoMME, MLVU, and Vinoground. TP and TR denotes Temporal Perception and Temporal Reasoning, respectively, two subtasks of VideoMME.
@@ 3.3 Analysis
   + TEMPO (2B-version data) 60.9 64.3 37.4 28.2 12.4 26.0
@@
   + TEMPO (7B-version data) 52.9 59.5 26.0 24.6 6.2 18.9
@@ 4 Related Work
Video Large Languge Models
@@
Direct Preference Optimization
Originally introduced in the text domain as an alternative to Reinforcement Learning from Human Feedback (RLHF) [23], Direct Preference Optimization (DPO) [24] was first employed to align language model behavior with human preferences, particularly to mitigate the generation of harmful or undesirable responses. DPO optimizes an implicit reward model derived from preference data, providing a straightforward and effective approach to post-training alignment. Due to its simplicity and strong performance, DPO has gained widespread adoption, leading to various improvements such as KTO [7] and $\beta$-DPO [32]. More recently, research has explored the adaptation of DPO and other alignment strategies to multi-modal LLMs by developing carefully curated data generation pipelines. For instance, POVID [41] constructs preference data for image-language models by comparing responses to clean and noised visual inputs, while MIA-DPO [20] generates preference data for multi-image understanding by concatenating single-image datasets and assessing whether model responses correctly attend to the intended image. Along with some concurrent work [38, 15, 34], we take a pioneering step toward leveraging DPO for improving the temporal reasoning capabilities of Video LLMs, addressing a key limitation that has been frequently highlighted in prior studies.

### v2

# TEMPLE: Temporal Preference Learning of Video LLMs via Difficulty Scheduling and Pre-SFT Alignment

@@ Abstract
Video Large Language Models (Video LLMs) have achieved significant success by leveraging a two-stage paradigm: pretraining on large-scale video-text data for vision-language alignment, followed by supervised fine-tuning (SFT) for task-specific capabilities. However, existing approaches struggle with temporal reasoning due to weak temporal correspondence in the data and reliance on the next-token prediction paradigm during training. To address these limitations, we propose TEMPLE (TEMporal Preference Learning), a systematic framework that enhances Video LLMs’ temporal reasoning capabilities through Direct Preference Optimization (DPO). To facilitate this, we introduce an automated preference data generation pipeline that systematically constructs preference pairs by selecting videos that are rich in temporal information, designing video-specific perturbation strategies, and finally evaluating model responses on clean and perturbed video inputs. Our temporal alignment features two key innovations: curriculum learning which that progressively increases perturbation difficulty to improve model robustness and adaptability; and “Pre-SFT Alignment”, applying preference optimization before instruction tuning to prioritize fine-grained temporal comprehension. Extensive experiments demonstrate that our approach consistently improves Video LLM performance across multiple benchmarks with a relatively small set of self-generated DPO data. We further analyze the transferability of DPO data across architectures and the role of difficulty scheduling in optimization. Our findings highlight our TEMPLE as a scalable and efficient complement to SFT-based methods, paving the way for developing reliable Video LLMs. Code is available at https://github.com/lscpku/TEMPLE.
@@ 1 Introduction
To address these limitations, in this paper, we propose TEMPLE (TEMporal Preference LEarning), a systematic framework that tackles both challenges through: (1) a scalable, automatic dataset curation pipeline that generates high-quality temporal preference data without relying on external LLMs, and (2) a tailored temporal alignment methodology that goes beyond next-token prediction. TEMPLE is built on Direct Preference Optimization (DPO) [24], which optimizes the relative quality of model responses by maximizing the reward margin between preferred and non-preferred outputs. Unlike SFT, which requires high-quality, human-verified responses, DPO focuses on relative comparisons rather than absolute correctness—making it particularly effective for video tasks where perfect annotations are difficult to obtain. Subtle temporal aspects like sequence order and motion speed are critical for comprehensive video understanding yet challenging to annotate manually at scale. Our preference-based framework naturally discourages shortcut learning while enabling efficient scaling through automated annotation, making TEMPLE an ideal methodology for enhancing temporal reasoning in Video LLMs.
@@
We implement TEMPLE through an automatic preference data generation pipeline that establishes the foundation for effective temporal alignment in Video LLMs. Inspired by advances in image-based LLMs [33, 41], we generate preference pairs by evaluating model responses under both clean and corrupted visual inputs. Our pipeline consists of three key stages: (1) curating a diverse set of videos that emphasize complex temporal relationships; (2) introducing targeted temporal perturbations—including frame dropping, sequence reversal, and frame shuffling—that precisely target known temporal reasoning weaknesses of current Video LLMs; and (3) constructing preference pairs by contrasting the model’s outputs for clean versus temporally corrupted videos. We also introduce a difficulty factor that quantifies perturbation magnitude to control the sample difficulties. This systematic approach enables us to efficiently generate high-quality contrastive examples that directly highlight temporal understanding capabilities without requiring expensive human annotation.
@@
Our tailored temporal alignment framework consists of two key technical innovations. First, we implement curriculum learning with progressive difficulty lifting, which gradually increases sample complexity during optimization. This controlled exposure to increasingly challenging temporal contrastive pairs enables more effective learning. Second, we introduce a novel “Pre-SFT Alignment” strategy, which reverses the conventional order of model alignment. Traditional approaches apply DPO after instruction-tuning (SFT) [30], assuming the model has already learned basic capabilities. In contrast, we argue that robust temporal understanding should be established first, before learning to follow diverse instructions. Our Pre-SFT Alignment applies DPO first on our curated dataset, forcing the model to develop strong temporal reasoning and video-text correspondence prior to any instruction tuning. These innovations, coupled with our carefully crafted self-sufficient video filtering and preference generation pipeline, distinguish our work from concurrent works like Tarsier 2 [34] and TPO [15], which rely on external LLMs (e.g., ChatGPT for Tariser 2 and GPT-4o for TPO) for data filtering and apply DPO in a post-SFT manner.
@@
Through extensive evaluation, we demonstrate that TEMPLE consistently improves model performance across multiple benchmarks including Video-MME [8], MLVU [40] and Vinoground [36] with only a small number of self-generated DPO pairs—eliminating dependence on external models or APIs. Our ablation studies validate the effectiveness of both technical contributions: the difficulty curriculum significantly stabilizes training and improves final performance, while the Pre-SFT Alignment strategy significantly outperforms the conventional SFT-then-DPO approach, particularly on tasks requiring fine-grained temporal reasoning. Further analysis reveals valuable insights about DPO data transferability between different model architectures and scales, demonstrating the framework’s flexibility. By introducing this automated, structured preference learning pipeline, we advance the state of video LLMs and pave the way for more robust and temporally-aware multi-modal models.
@@ 3 Experiment
   + TEMPLE (Ours) 7B 74.5  (+3.6) 45.2  (+2.8) 61.0  (+0.6) 65.8  (+1.3) 37.2  (+3.2) 29.6  (+3.0) 12.2  (+1.6) 26.3  (+2.6)
Table 1: Benchmark results of our approach on VideoMME, MLVU, and Vinoground. TP and TR denote Temporal Perception and Temporal Reasoning, respectively, two subtasks of VideoMME.
@@ 3.3 Analysis
   + TEMPLE (2B-version data) 60.9 64.3 37.4 28.2 12.4 26.0
@@
   + TEMPLE (7B-version data) 52.9 59.5 26.0 24.6 6.2 18.9
@@ 4 Related Work
4.1 Video Large Languge Models
@@
4.2 Direct Preference Optimization
Originally introduced in the text domain as an alternative to Reinforcement Learning from Human Feedback (RLHF) [23], Direct Preference Optimization (DPO) [24] was first employed to align language model behavior with human preferences, particularly to mitigate the generation of harmful or undesirable responses. DPO optimizes an implicit reward model derived from preference data, providing a straightforward and effective approach to post-training alignment. Due to its simplicity and strong performance, DPO has gained widespread adoption, leading to various improvements such as KTO [7] and $\beta$-DPO [32]. More recently, research has explored the adaptation of DPO and other alignment strategies to multi-modal LLMs by developing carefully curated data generation pipelines. For instance, POVID [41] constructs preference data for image-language models by comparing responses to clean and noised visual inputs, while MIA-DPO [20] generates preference data for multi-image understanding by concatenating single-image datasets and assessing whether model responses correctly attend to the intended image.
Recent concurrent works [38, 15, 34] have also explored DPO for video LLMs based on model responses to corrupt visual inputs, boosting model performance on a wide range of tasks. However, our approach distinguishes itself through several key innovations. For instance, Tarsier 2 [34] applies DPO exclusively to short video clips, limiting its ability to capture long-range temporal dependencies. Additionally, both Tarsier 2 and TPO [15] are reliant on external proprietary models to curate and filter preference data. In contrast, our method introduces a carefully designed pipeline that automatically filters and selects temporally rich videos spanning 1–3 minutes and generates preference data in a fully self-sufficient manner. Furthermore, beyond dataset curation, we present two important technical innovations: difficulty scheduling, which structures training based on complexity for more effective learning, and pre-SFT alignment, which injects temporal knowledge before instruction tuning. These innovations collectively contribute to more robust and temporally-aware video reasoning, setting our approach apart from concurrent work in video DPO.